OpenStudy (anonymous):

suppose a parabola has an axis of symmetry at x=-7, a maximum height of 4 and also passes through the point (-6,0). Write the parabola in vertex form.

OpenStudy (nikato):

Do u know what the vertex form of a parabola is?

OpenStudy (anonymous):

no clue. im thinking the answer was y=a(x-h^2)+k where a= -7, h= 2/7, and k= -38/7

OpenStudy (nikato):

No, the numbers r incorrect. And btw, thats^ called the vertex form of a parabola equation

OpenStudy (campbell_st):

well you know the vertex.... its x = -7 and y = 4 the line of symmetry and max value for y are the vertex so using your formula h = -7 and k = 4 substitute the point (-6, 0) to find the value of a. hope this helps

OpenStudy (campbell_st):

oh, the parabola is concave down... because the question mentions maximum... y = 4

OpenStudy (anonymous):

thanks guys. i appreciate your help!